If a triangle has a base of 2x+1 and a height of 6x-3. What value of X would give and Area of 240 m^2.
Please help me on how to put this into an equation and solve it.

Answer by nerdybill(7384)

If a triangle has a base of 2x+1 and a height of 6x-3. What value of X would give and Area of 240 m^2.
.
Area of any triangle is (1/2)bh
where
b is base
h is height
.
In your case then:
(1/2)(2x+1)(6x-3) = 240
multiplying both sides by 2:
(2x+1)(6x-3) = 480
FOIL the left side:
12x^2 - 6x + 6x - 3 = 480
12x^2 - 3 = 480
12x^2 = 483
x^2 = 483/12
x = 6.3443